# Communication-Efficient Federated Optimization over Semi-Decentralized Networks

### v2

@@ 3 Proposed Algorithm
5:     for $t=1,2\cdots,T_{o}$ donormal-$\triangleright$ $T_{o}$-local updates
6:         Draw the mini-batch $\mathcal{Z}^{k+1,t}=\{\mathcal{Z}_{i}^{k+1,t}\}_{i=1}^{n}$ and compute
$\boldsymbol{X}^{k+1,t}$ $=\boldsymbol{X}^{k+1,t-1}-\eta_{l}\boldsymbol{Y}^{k+1,t-1},$ (3a)
$\boldsymbol{G}^{k+1,t}$ $=\frac{1}{b}\nabla\ell(\boldsymbol{X}^{k+1,t};\mathcal{Z}^{k+1,t}),$ (3b)
@@
8:     Define $\boldsymbol{W}^{k}=\begin{cases}\boldsymbol{J}&\text{with probability}~{}p,\\ \boldsymbol{W}&\text{otherwise}.\end{cases}$$\triangleright$ probablistic heterogeneous communication
9:     Draw the mini-batch $\mathcal{Z}^{k+1}=\{\mathcal{Z}_{i}^{k+1}\}_{i=1}^{n}$ and update
$\boldsymbol{X}^{k+1}$ $=\left((1-\eta_{c})\boldsymbol{X}^{k}+\eta_{c}(\boldsymbol{X}^{k+1,T_{o}}-\eta_{l}\boldsymbol{Y}^{k+1,T_{o}})\right)\boldsymbol{W}^{k},$ (4a)
$\boldsymbol{G}^{k+1}$ $=\frac{1}{b}\nabla\ell(\boldsymbol{X}^{k+1};\mathcal{Z}^{k+1}),$ (4b)
@@
  • The second stage is to perform information exchange over the semi-decentralized network via a probabilistic communication model (cf. line 8-10), i.e., there are two possible communication schemes — agent-to-server communication with probability $p$ and agent-to-agent communication otherwise. Different schemes correspond to different mixing matrices (cf. line 8), i.e., if agents implement the global communication, set $\boldsymbol{W}^{k}=\boldsymbol{J}$; otherwise, set $\boldsymbol{W}^{k}=\boldsymbol{W}$. Then, agents update $\{\boldsymbol{X}^{k+1},\boldsymbol{Y}^{k+1},\boldsymbol{G}^{k+1}\}$ via (9), using the output of local updates $\{\boldsymbol{X}^{k+1,T_{o}},\boldsymbol{Y}^{k+1,T_{o}},\boldsymbol{G}^{k+1,T_{o}}\}$ via the selected communication scheme.
@@ 4.1 Assumptions
Given the underlying graph $\mathcal{G}$ and its mixing matrix $\boldsymbol{W}$ following the Definition 1, then $\boldsymbol{W}^{k}$ defined in Algorithm 1 satisfies that for any $\boldsymbol{x}\in\mathbb{R}^{n}$
$\mathbb{E}[\|\boldsymbol{W}^{k}\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2}]\leq(1-\lambda_{p})\|\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2},$
@@
Noted that Assumption 1 is weaker than the connected assumption in prior semi-decentralized literatures [${}^{+}$21, GSHG22, WFC22], i.e., $\lambda_{w}>0$. More specifically, Assumption 1 implies that the underlying graph can be disconnected if and only if $p>0$. Only in the case that the centralized server is unavailable (i.e., $p=0$), Assumption 1 presumes the connectivity of $\mathcal{G}$, i.e., $\lambda_{p}=\lambda_{w}>0$.
Regarding to the objective functions, we assume that the optimal value $f^{\star}\coloneqq\min_{\boldsymbol{x}}f(\boldsymbol{x})$ exists and $f^{\star}>-\infty$. The local objective functions $\{f_{i}\}_{i=1}^{n}$ could be nonconvex but satisfy the standard smoothness assumption provided below.
@@
Each local function $f_{i}(\boldsymbol{x})$ is differentiable and there exists a constant $L$ such that for each $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$:
$\|\nabla f_{i}(\boldsymbol{x})-\nabla f_{i}(\boldsymbol{y})\|_{2}\leq L\|\boldsymbol{x}-\boldsymbol{y}\|_{2}.$
In addition, we assume that the local stochastic gradient $\boldsymbol{g}_{i}$ is an unbiased estimate with a bounded variance, which is widely used in the literature [${}^{+}$21, WFC22, GSHG22, GC23, KLS21, ${}^{+}$].
Assumption 3 (Unbiased gradient and bounded variance).
For every agent $i\in[n]$, the local stochastic gradient $\boldsymbol{g}_{i}$ is unbiased, i.e., $\mathbb{E}[\boldsymbol{g}_{i}]=\nabla f_{i}(\boldsymbol{x}_{i})$, and there exists a constant $\sigma\geq 0$ s.t.
$\mathbb{E}_{\mathcal{Z}_{i}\sim\mathcal{D}_{i}}[\|\boldsymbol{g}_{i}-\nabla f_{i}(\boldsymbol{x}_{i})\|_{2}^{2}]\leq\sigma^{2}/b.$
or in the compact form
$\mathbb{E}_{\mathcal{Z}\sim\mathcal{D}}[\|\boldsymbol{G}-\nabla F(\boldsymbol{X})\|_{\mathsf{F}}^{2}]\leq\frac{n\sigma^{2}}{b}.$
Note that in the case of the full-batch gradients, i.e., the mini-batch size $b=m$, we can simply set $\sigma=0$ and thus Assumption 3 still holds.
@@ 4.2 Convergence analysis of PISCO
Now, we are ready to present our main results. First, the following theorem demonstrates that our proposed PISCO algorithm is able to converge to the neighborhood of the stationary solution to the problem (1) at a rate of $O(1/K)$ with constant step-sizes.
Theorem 1 (Convergence rate with mini-batch gradients).
Suppose Assumption 1, 2 and 3 hold. Let $\tilde{f}=f(\overline{\boldsymbol{x}}^{0})-f^{\star}$ and $\boldsymbol{\Phi}_{y}^{0}=\boldsymbol{Y}^{0}-\boldsymbol{Y}^{0}\boldsymbol{J}$. For any $\alpha\geq 1$ s.t. $\eta_{c}=\alpha\sqrt{(1+p)}\lambda_{p},\eta_{l}\leq\frac{\sqrt{(1+p)}\lambda_{p}}{360\alpha L(T_{o}+1)}$, we have
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ (5)
@@
Proof.
See Appendix A. ∎
Noted that the above convergence result can hold even under significant data heterogeneity across agents, since we do not assume any bounded similarity between local objectives. Because of the existence of local variance $\sigma^{2}$, we fine-tune the local-update step-size to obtain the exact convergence rate with the leading term $O(1/\sqrt{nT_{o}K})$, given by the following corollary.
Corollary 1 (Convergence rate with mini-batch gradients).
Suppose all the conditions in Theorem 1 hold. If the step-sizes satisfy $\eta_{c}=\alpha\sqrt{1+p}\lambda_{p}$ and
$\eta_{l}=\frac{1}{\alpha T_{o}}\min\{\left(\frac{nbT_{o}\tilde{f}}{(1+p)\lambda_{p}^{2}L\sigma^{2}K}\right)^{1/2},\left(\frac{\sqrt{1+p}\lambda_{p}b\tilde{f}}{(1-p)L^{2}\sigma^{2}K}\right)^{1/3},\left(\frac{\alpha^{2}b\tilde{f}}{\sqrt{1+p}\lambda_{p}L^{2}\sigma^{2}K}\right)^{1/3}\}$
and $K$ is sufficiently large to satisfy (47), PISCO converges at the rate of
$O\left(\left(\frac{L\sigma^{2}}{nT_{o}bK}\right)^{\frac{1}{2}}+\left(\frac{L\sigma}{\alpha\sqrt{1+p}\lambda_{p}\sqrt{b}K}\right)^{\frac{2}{3}}+\left(\frac{\sqrt{1-p}L\sigma}{(1+p)\lambda_{p}^{2}\sqrt{b}K}\right)^{\frac{2}{3}}+\frac{1}{nK}\right).$ (6)
Proof.
See Appendix B. ∎
@@
$O\left(\frac{L\sigma^{2}}{nT_{o}b\epsilon^{4}}+\frac{L\sigma}{\alpha\sqrt{1+p}\lambda_{p}\sqrt{b}\epsilon^{3}}+\frac{\sqrt{1-p}L\sigma}{(1+p)\lambda_{p}^{2}\sqrt{b}\epsilon^{3}}+\frac{1}{n\epsilon^{2}}\right)$ (7)
communication rounds. Notice that if $K$ is sufficiently large and the first term $\left(\frac{L\sigma^{2}}{nT_{o}bK}\right)^{\frac{1}{2}}$ correspondingly becomes dominant, increasing the number of agents $n$ or the number of local updates $T_{o}$ can accelerate the convergence. Such a speedup matches the findings in the special cases of semi-decentralized ML, i.e., decentralized setting [LLKS23, GC23] when $p=0$ and federated setting [${}^{+}$20] when $p=1$.
@@
Remark 1 (Decentralized cases).
which is better than the rate of Periodical-GT in [LLKS23] and LSGT [GC23], since the network dependency is $O(1/\lambda_{w}^{2})$ and only appears in the second term (see Table 2). The second term is slightly worse than $K$-GT, the variance-reduction version of Periodical-GT [LLKS23], because it corrects the descent direction as the average of $T_{o}$ local updates instead of the last local update at communication. However, their theoretical guarantee requires the initial local correction variables to be settled in a centralized way. Combining our analysis with such a variance-reduction method presented in [LLKS23] while avoiding the centralized initialization would be a promising future direction of this paper.
Remark 2 (Federated case).
When $p=1$, every agent can communicate with the server directly and thus PISCO performs like federated algorithms with local updates at any iteration $k\geq 1$. Correspondingly, the communication complexity becomes
where the leading term is the same as that of SCAFFOLD in [${}^{+}$20] with the linear speedup in terms of the network size $n$ and the number of local updates $T_{o}$.
Moreover, the highlight of our work is to fill the void of semi-decentralized ML with the probabilistic communication model and gain the best communication efficiency from both agent-to-agent communication and agent-to-server communication, as shown in Table 2. In addition, PISCO is able to improve the network dependency of the communication overheads from $O(\lambda_{w}^{-2})$ [LLKS23] to $O(\lambda_{p}^{-2})$, where the trade-off between the communication probability and the network dependency is illustrated in Figure 3. The flexible heterogeneous communication brings the superior communication efficiency of PISCO in both well-connected and poorly-connected networks.
@@
Table 2: The number of (expected) server/local communication rounds of existing distributed algorithms with stochastic gradients and local updates, to achieve the $\epsilon$-accuracy. Here, $n$ is the number of agents, $T_{o}$ is the number of local updates, $\lambda_{w}$ is the mixing rate of the underlying graph.
Figure 3: The network dependency of PISCO regarding to agent-to-server communication probability $p$.
Implications for well-connected networks.
@@
Implications for poorly-connected networks.
When $\lambda_{w}\to 0$, performing agent-to-agent only communication often results in a large number of communication rounds and prohibitive communication costs. Surprisingly, as shown in Figure 3, with any probability $p\geq\lambda_{w}$, the network dependency can be reduced to $O(p^{-2})$. More specifically, even a small agent-to-server probability $p=O(\sqrt{\lambda_{w}})$ can significantly improve the network dependency from $O(\lambda_{w}^{-2})$ to $O(\lambda_{w}^{-1})$. Take the large-scale path graph as an example, where the mixing rate $\lambda_{w}$ scales on the order of $O(1/n^{2})$ [NOR18]. Our PISCO with $p=O(1/n)$ can improve the network dependency from $O(n^{4})$ to $O(n^{2})$. Moreover, if $p=O(1)$, the communication complexity can be network-independent like Gossip-PGA [${}^{+}$21], but our theoretical analysis does not require the additional assumption of the bounded dissimilarity between local objectives.
In many real-world scenarios, it is also popular to choose large mini-batch size $b$ to guarantee the exact convergence to the stationary point. As the terms related to local variance on the RHS of (5) scale on the order of $O(\sigma^{2}/b)$, if we choose a large enough mini-batch size $b=O(\sigma^{2}/\epsilon^{2})$, the following result holds.
Corollary 2 (Communication complexity with large batch size).
Suppose all the conditions in Theorem 1 holds. If the mini-batch size is sufficiently large, i.e., $O(\frac{\sigma^{2}}{\epsilon^{2}})$, it holds $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\bar{x}^{k})\|_{2}^{2}]\leq\epsilon^{2}$ after
@@
Note that Corollary 2 matches the result in decentralized setting [NAYU22] and federated setting [${}^{+}$20], by setting $p=0$ and $p=1$ respectively.

### v4

@@ 3 Proposed PISCO Algorithm
5:     for $t=1,2\cdots,T_{o}$ do
6:         Draw the mini-batch $\mathcal{Z}^{k+1,t}$ and compute
$\boldsymbol{X}^{k+1,t}$ $=\boldsymbol{X}^{k+1,t-1}-\eta_{l}\boldsymbol{Y}^{k+1,t-1}$ (3a)
$\boldsymbol{G}^{k+1,t}$ $=\frac{1}{b}\nabla\ell(\boldsymbol{X}^{k+1,t};\mathcal{Z}^{k+1,t})$ (3b)
@@
8:     Define $\boldsymbol{W}^{k}=\begin{cases}\boldsymbol{J}&\text{with probability}~{}p,\\ \boldsymbol{W}&\text{otherwise}.\end{cases}$
9:     Draw the mini-batch $\mathcal{Z}^{k+1}$ and update
$\boldsymbol{X}^{k+1}$ $=\left((1-\eta_{c})\boldsymbol{X}^{k}\!+\!\eta_{c}(\boldsymbol{X}^{k+1,T_{o}}-\eta_{l}\boldsymbol{Y}^{k+1,T_{o}})\right)\boldsymbol{W}^{k}$ (4a)
$\boldsymbol{G}^{k+1}$ $=\frac{1}{b}\nabla\ell(\boldsymbol{X}^{k+1};\mathcal{Z}^{k+1})$ (4b)
@@
  • The second stage is to perform the information exchange over the semi-decentralized network via a probabilistic communication model (cf. line 8-10), i.e., there are two possible communication schemes — agent-to-server communication with probability $p$ and agent-to-agent communication otherwise. Different schemes correspond to different mixing matrices (cf. line 8), i.e., if agents implement the global communication, set $\boldsymbol{W}^{k}=\boldsymbol{J}$; otherwise, set $\boldsymbol{W}^{k}=\boldsymbol{W}$. Then, agents update $\{\boldsymbol{X}^{k+1},\boldsymbol{Y}^{k+1},\boldsymbol{G}^{k+1}\}$ via (4), using the output of local updates $\{\boldsymbol{X}^{k+1,T_{o}},\boldsymbol{Y}^{k+1,T_{o}},\boldsymbol{G}^{k+1,T_{o}}\}$ via the selected communication scheme.
@@ 4.1 Assumptions
Given the undirected graph $\mathcal{G}$ and its mixing matrix $\boldsymbol{W}$ following the Definition 1, then $\boldsymbol{W}^{k}$ defined in Algorithm 1 satisfies
$\mathbb{E}[\|\boldsymbol{W}^{k}\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2}]\leq(1-\lambda_{p})\|\boldsymbol{x}-\overline{\boldsymbol{x}}\|_{2}^{2},\quad\forall\boldsymbol{x}\in\mathbb{R}^{n},$
@@
Note that Assumption 1 is weaker than the connected assumption in prior semi-decentralized literatures [CYZ+21, GSHG22, WFC22], i.e., $\lambda_{w}>0$. More specifically, Assumption 1 implies that the underlying graph can be disconnected if and only if $p>0$. Only in the case that the centralized server is unavailable (i.e., $p=0$), Assumption 1 presumes the connectivity of $\mathcal{G}$.
Regarding the objective functions, we assume that the optimal value $f^{\star}\coloneqq\min_{\boldsymbol{x}}f(\boldsymbol{x})$ exists and $f^{\star}>-\infty$. The local objective functions $\{f_{i}\}_{i=1}^{n}$ could be nonconvex but satisfy the standard smoothness assumption provided below.
@@
Each local function $f_{i}(\boldsymbol{x})$ is differentiable and there exists a constant $L$ such that
$\|\nabla f_{i}(\boldsymbol{x})-\nabla f_{i}(\boldsymbol{y})\|_{2}\leq L\|\boldsymbol{x}-\boldsymbol{y}\|_{2},\quad\forall\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}.$
In addition, we assume that the local stochastic gradient $\boldsymbol{g}_{i}$ is an unbiased estimate with a bounded variance, which is widely used in the literature [CYZ+21, WFC22, GSHG22, GC23, KLS21, ZLL+22].
Assumption 3 (Bounded variance).
For every agent $i\in[n]$, there exists a constant $\sigma\geq 0$ s.t.
$\mathbb{E}_{\mathcal{Z}_{i}\sim\mathcal{D}_{i}}[\|\boldsymbol{g}_{i}-\nabla f_{i}(\boldsymbol{x})\|_{2}^{2}]\leq\sigma^{2}/b,\quad\forall\boldsymbol{x}\in\mathbb{R}^{d}.$
Note that in the case of the full-batch gradients, i.e., the mini-batch size $b=m$, we can simply set $\sigma=0$ and thus Assumption 3 always holds.
@@ 4.2 Convergence analysis of PISCO
Now, we are ready to present our main results. First, the following theorem demonstrates that our proposed PISCO  is able to converge to the neighborhood of the stationary solution to the problem (1) at the rate of $O(1/K)$ with constant step-sizes. The proof is postponed to the Appendix B.
Theorem 1 (Convergence rate).
Suppose Assumption 1, 2 and 3 hold. Let $\tilde{f}=f(\overline{\boldsymbol{x}}^{0})-f^{\star}$ and $\boldsymbol{\Phi}_{y}^{0}=\boldsymbol{Y}^{0}-\boldsymbol{Y}^{0}\boldsymbol{J}$. For any $\alpha\geq 0.1$ s.t. $\eta_{c}=\alpha\sqrt{(1+p)}\lambda_{p}$ and $\eta_{l}\leq\frac{\sqrt{(1+p)}\lambda_{p}}{360\alpha L(T_{o}+1)}$, it holds that $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ converges at the rate of
@@
Note that the above convergence result can hold even under significant data heterogeneity across agents, since we do not assume any bounded similarity between local objectives.
Due to the existence of the variance $\sigma^{2}$, we fine-tune the local-update step-size to obtain the exact convergence rate with the leading term $O(1/\sqrt{nT_{o}K})$, based on Theorem 1. Specifically, the following corollary considers the scenarios with mini-batch gradients (i.e., $b\leq O(\sigma^{2}K)$ and $\sigma>0$), while the case of large or full batch gradients (i.e., the batch size $b\geq\Theta(\sigma^{2}K)$ or $\sigma=0$) will be discussed later in Corollary 2. The proof of Corollary 1 is postponed to Appendix C.
Corollary 1 (Convergence rate with mini batch).
Suppose all the conditions in Theorem 1 hold. Consider that the number of communication rounds $K$ is sufficiently large, i.e., $K\geq\Theta(\frac{nbT_{o}L\tilde{f}}{\lambda_{p}^{4}\sigma^{2}})$, and the mini-batch size $b\leq\!O(\sigma^{2}K)$, where $\sigma>\!0$. If the step-sizes $\eta_{c}=\alpha\sqrt{1+p}\lambda_{p}$, $\eta_{l}=\frac{1}{\alpha T_{o}}\min\big{\{}\sqrt{\frac{n\alpha^{2}bT_{o}\tilde{f}}{\eta_{c}^{2}L\sigma^{2}K}},\sqrt[3]{\frac{\eta_{c}b\tilde{f}}{\alpha L^{2}\sigma^{2}K}}\big{\}}$. Then, $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\overline{\boldsymbol{x}}^{k})\|_{2}^{2}]$ can converge at the rate of
$O\left(\left(\frac{L\sigma^{2}}{nT_{o}bK}\right)^{\frac{1}{2}}\!\!+\left(\frac{L\sigma}{\lambda_{p}^{2}\sqrt{b}K}\right)^{\frac{2}{3}}\!\!+\!\frac{1}{nK}\right).$
@@
$O\left(\frac{L\sigma^{2}}{nT_{o}b\epsilon^{4}}+\frac{L\sigma}{\lambda_{p}^{2}\sqrt{b}\epsilon^{3}}+\frac{1}{n\epsilon^{2}}\right)$
communication rounds. Notice that if $K$ is sufficiently large and the first term $\left(\frac{L\sigma^{2}}{nT_{o}bK}\right)^{\frac{1}{2}}$ correspondingly becomes dominant, increasing the number of agents $n$ or the number of local updates $T_{o}$ can accelerate the convergence. Such a linear speedup matches the findings in the special cases of semi-decentralized ML, i.e., decentralized setting [LLKS24, GC23] when $p=0$ and federated setting [KKM+20] when $p=1$.
@@
Table 2: The number of the expected agent-to-server/agent communication rounds of ours and existing decentralized and federated ML algorithms with stochastic gradients and local updates, to achieve the $\epsilon$-accuracy, where $\epsilon$ is sufficiently small. Here, $n$ is the number of agents, $T_{o}$ is the number of local updates, $\lambda_{w}$ is the mixing rate of the underlying graph.
Figure 3: The network dependency of PISCO regarding agent-to-server communication probability $p$.
Remark 1 (Decentralized case).
which is better than the rate of Periodical-GT in [LLKS24] and LSGT [GC23], since the network dependency is $O(1/\lambda_{w}^{2})$ and only appears in the second term (see Table 2). The second term is slightly worse than $K$-GT, the variance-reduced Periodical-GT [LLKS24], since it corrects the descent direction with the average of $T_{o}$ local updates instead of the last local update at communication. However, they require that the initial local correction variables are settled in a centralized way. Combining our analysis with such a variance-reduction method while avoiding the centralized initialization would be a promising future direction of this paper.
Remark 2 (Federated case).
When $p=1$, every agent can communicate with the server directly and thus PISCO performs in the federated fashion at any iteration $k\geq 1$. Then, the communication complexity becomes
where the leading term is the same as that of SCAFFOLD in [KKM+20] with the linear speedup in terms of the network size $n$ and the number of local updates $T_{o}$.
Moreover, the highlight of our work is to fill the void of semi-decentralized ML with the probabilistic communication model and gain the best communication efficiency from both agent-to-agent communication and agent-to-server communication, as shown in Table 2. In addition, PISCO is able to improve the network dependency of the communication overheads from $O(\lambda_{w}^{-2})$ [LLKS24] to $O(\lambda_{p}^{-2})$, where the trade-off between the communication probability and the network dependency is illustrated in Figure 3. The flexible heterogeneous communication brings the superior communication efficiency of PISCO in both well-connected and poorly-connected networks.
Remark 3 (For well-connected networks).
@@
Remark 4 (For poorly-connected networks).
When $\lambda_{w}\to 0$, performing agent-to-agent only communication often results in a large number of communication rounds and prohibitive communication costs. As shown in Figure 3, with any probability $p\geq\lambda_{w}$, the network dependency can be reduced to $O(p^{-2})$. More specifically, even a small agent-to-server probability $p=\Theta(\sqrt{\lambda_{w}})$ can significantly improve the network dependency from $O(\lambda_{w}^{-2})$ to $O(\lambda_{w}^{-1})$. Take the large-scale path graph as an example, where the mixing rate $\lambda_{w}$ scales on the order of $O(1/n^{2})$ [NOR18]. Our PISCO with $p=\Theta(1/n)$ can improve the network dependency from $O(n^{4})$ to $O(n^{2})$. Moreover, if $p=\Theta(1)$, the communication complexity can be network-independent like Gossip-PGA [CYZ+21], but our theoretical analysis does not require the additional assumption of the bounded dissimilarity between local objectives.
In many real-world scenarios, it is also popular to choose large mini-batch size $b$ to guarantee the exact convergence to the stationary point. As the terms related to the variance $\sigma^{2}$ on the right hand side of (5) scale on the order of $O(\sigma^{2}/b)$, if we choose a large enough mini-batch size $b\geq\Theta(\sigma^{2}/\epsilon^{2})$, the following desirable result holds.
Corollary 2 (Communication complexity with large batch).
Suppose all the conditions in Theorem 1 holds. If the batch size $b$ is sufficiently large, i.e., $b\geq\Theta(\frac{\sigma^{2}}{\epsilon^{2}})$, it holds $\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[\|\nabla f(\bar{\boldsymbol{x}}^{k})\|_{2}^{2}]\leq\epsilon^{2}$ after
@@
Note that Corollary 2 also matches the result in decentralized setting [NAYU23] and federated setting [KKM+20], by setting $p=0$ and $p=1$ respectively.